You observe that the U.K. annual interest rate is 2.5 percent, the U.S. annual interest rate is 3.8 percent, the 3-month forward rate is $1.8180/E, and the spot rate is $1.8034/t. Assuming that transaction costs are 0.2 percent, are the financial markets in equilibrium?
Answer #1

Solution:

Computation of quarterly interest rate differential: {(0.038 - 0.025) / 4} = 0.00325

Computation of Forward premium: {(1.8180-1.8034)/(1.8034)} = 0.00896

Forward premium after transactions costs: (0.00896 - 0.002) = 0.00696

Thus we can say that it would be profitable to invest in the foreign market